# GLINT-RU: Gated Lightweight Intelligent Recurrent Units for Sequential Recommender Systems

### v1

@@ Abstract.
In the rapidly evolving field of artificial intelligence, transformer-based models have gained significant attention in the context of Sequential Recommender Systems (SRSs), demonstrating remarkable proficiency in capturing user-item interactions. However, such attention-based frameworks result in substantial computational overhead and extended inference time. To address this problem, this paper proposes a novel efficient sequential recommendation framework GLINT-RU that leverages dense selective Gated Recurrent Units (GRU) module to accelerate the inference speed, which is a pioneering work to further exploit the potential of efficient GRU modules in SRSs. The GRU module lies at the heart of GLINT-RU, playing a crucial role in substantially reducing both inference time and GPU memory usage. Through the integration of a dense selective gate, our framework adeptly captures both long-term and short-term item dependencies, enabling the adaptive generation of item scores. GLINT-RU further integrates a mixing block, enriching it with global user-item interaction information to bolster recommendation quality. Moreover, we design a gated Multi-layer Perceptron (MLP) for our framework where the information is deeply filtered. Extensive experiments on three datasets are conducted to highlight the effectiveness and efficiency of GLINT-RU. Our GLINT-RU achieves exceptional inference speed and prediction accuracy, outperforming existing baselines based on Recurrent Neural Network (RNN), Transformer, MLP and State Space Model (SSM). These results establish a new standard in sequential recommendation, highlighting the potential of GLINT-RU as a renewing approach in the realm of recommender systems. The implementation code is available at https://github.com/szhang-cityu/GLINT-RU.
@@ 1. INTRODUCTION
In this era of data exploding, Sequential Recommender Systems (SRSs) (Zhang et al., 2019; Yue et al., 2024; Liu et al., 2023a; Hidasi et al., 2015; Kang and McAuley, 2018; Liu et al., 2024; Li et al., 2022) have gained much attention in capturing users’ preferences within a large amount of sequential interaction data. GRU4Rec (Hidasi et al., 2015) as one of the earliest session-based recommendation models, employs stacked Gated Recurrent Units (GRU) for item-to-item recommendations. However, the RNN-based methods (Hidasi et al., 2015; Li et al., 2017) are fading from the recommendation realm due to their relatively lower accuracy. In recent years, transformer-based SRSs (Kang and McAuley, 2018; Sun et al., 2019) have become increasingly popular for the powerful multi-head attention mechanism (Vaswani et al., 2017). They exhibit remarkable ability in capturing sequential interactions and delivering accurate predictions (Li et al., 2023). However, despite their effectiveness, current attention-based models suffer from substantial computational demands and extended inference time, which is caused by the dot product operation in the attention mechanism (Wang et al., 2020; Liu et al., 2023a).
@@
To tackle the issue of the high computational cost of attention-based SRSs, much research has been done to improve the attention mechanism to the linear computational complexity. LinRec (Liu et al., 2023a) changes the order of dot product between query and key matrices by designing a special mapping, dramatically reducing the inference time. LightSAN (Fan et al., 2021) projects historical interactions into interest representations with shorter length, thereby reducing the computational complexity of transformers to a linear scale. Mamba4Rec (Liu et al., 2024) outperforms the existing attention mechanism by utilizing one efficient mamba block based on a selective SSM variant (i.e., Mamba (Gu and Dao, 2023)), within which a structured state tensor is used to address long-range dependencies. SMLP4Rec (Gao et al., 2024) is an efficient pure Multi-layer Perceptron (MLP) framework that achieves fast inference by reshaping input sequence tensors, demonstrating the potential of MLP-based models (Li et al., 2022; Zhou et al., 2022; Long et al., 2024) for sequential recommendation tasks.
However, to achieve high performance, the transformer-based recommendation requires deeply stacked transformers. Even with a linearized attention mechanism, the inference speed of such large transformer architecture may not meet the desired efficiency standards (Zhang et al., 2024). Furthermore, the efficiency of the MLP architecture is adversely affected by the reshaping of large tensors, which is notably time-consuming. GRU4Rec (Hidasi et al., 2015) is also an efficient framework for recommendation tasks, but it exhibits limitations in learning complex user behaviors. In this paper we aim to further improve the efficiency of recommendations, and tackle the issue of low accuracy of traditional RNN-based frameworks.
@@
To further reduce resource consumption, accelerate inference speed, and enhance the model performance, we propose a novel efficient SRS framework named Gated Lightweight IntelligeNT Recurrent Units (GLINT-RU). This paper employs various gate mechanisms in appropriate positions to fully perceive the data environment and filter or select information automatically (Zhang et al., 2024; Chang et al., 2023; De et al., 2024). We introduce an expert mixing block that captures the long-/short-term dependencies via GRU, and utilize linear attention to capture the global interaction information between users and items (Jacobs et al., 1991; Masoudnia and Ebrahimpour, 2014). This strategy not only supplements the sequential information, but also improves the inference speed due to the linear computational complexity of paralleled GRU and linear attention mechanism. Additionally, we implement a dense selective GRU, which selects the output of GRU adaptively, and considers the connections among adjacent items. It leverages the gate mechanism to select crossed features, and extracts short-duration patterns to refine the model’s understanding of user behavior dynamics. Moreover, a gated MLP block is utilized to select on the outputs of the expert mixing block, deeply filtering the information based on the data environment.
We summarize the major contributions of our work as:
@@
In this paper, we introduce GLINT-RU, a novel and efficient recommendation framework that achieves remarkable inference speed with a streamlined architecture that only requires a single layer, and exhibits state-of-the-art performance. It is a advanced model that captures complex user-item interactions from mixing blocks, which substantially improves the performance of traditional sequential recommendation models.
@@
We introduce a dense selective GRU module, which not only incorporates connections between adjacent items, but also empowers the model with the capability to selectively learn sequential information. The integration of this advanced GRU module into the model markedly elevates its performance, establishing a new standard for efficient recommender systems.
We implement a gated MLP block, allowing for deep filtering of dense information by automatically perceiving the data environment. This mechanism enables our model to be more flexible and adaptive to complex sequential user behaviors.
@@ 2.1. Problem Statement
For a sequential recommendation task, we have a set of users who have historical interactions with a set of items . Among these users, the -th user has a preferred item sequence denoted as , , where is the length of the item list that the -th user interacts with. Our goal is to design an efficient framework and predict the next item that the user is most likely to interact.
@@ 2.2. Gated Recurrent Units
As an essential part of GLINT-RU framework, the GRU (Chung et al., 2014) module contributes to the recommendation task by capturing the long-term dependencies among the items in the sequence and dynamically adjust its memory content across different parts of the sequence (Dey and Salem, 2017; Shen et al., 2018). The update mechanism of a GRU cell can be formalized as follows:
@@
The GRU (Gated Recurrent Unit) module, equipped with update and reset gates in sequential GRU Cells, is adept at capturing the relationships among the items throughout the sequence, while maintaining a relatively low computational complexity. However, the sequential information in GRU cannot be interacted, and each hidden state is primarily encoded from preceding elements, which restricts the representational capacity of GRU-based recommendation models, particularly in capturing complex item dependencies across the entire sequence.
@@ 2.3. Linear Attention Mechanism
Attention mechanism as a powerful core of the transformer structure, exhibits performance in learning sequence interactions in recommendation tasks. However, the high computational cost of dot product between query matrix and key matrix substantially lower the inference speed of transformer-based SRSs especially when the sequence length is much larger than hidden size (Shen et al., 2021; Wang et al., 2020). To tackle this issue, the linear attention mechanism (Liu et al., 2023a) designs an special mapping function to change the order of the dot product and reduce the computational complexity to . The linear attention mechanism can be written as:
where and are row-wise and column-wise normalization mappings, respectively, are learnable query, key and value matrices and is the output attention score. This approach mitigates the issue that the softmax layer concentrates on the scores of merely a few positions, enlarging the information capacity of attention mechanism (Liu et al., 2023a). By implementing linear attention, our GLINT-RU framework is capable of learning interactions between items in long sequences.
@@ 3.1. Framework Overview
Many existing recommendation frameworks depend on transformer structure (Sun et al., 2019; Kang and McAuley, 2018; Liu et al., 2023a), which incurs substantial computational overhead and low inference speed. Restricted by the large computational complexity of stacked transformers, linear attention-based models have approached a plateau in terms of minimizing inference time and resource consumption. Uniquely, we propose an advanced recommendation framework that integrates the linear attention mechanism and efficient dense selective GRU module, which further reduce the computational cost compared with stacked linear transformers and SSM-based models. Additionally, this dense selective GRU module also enables our framework to understand complex long-/short-term user behaviors, and substantially reduce the computational cost and inference time. Figure 1 shows the structure for our GLINT-RU framework.
GLINT-RU framework integrates an expert mixing block for mixing sequential information from the dense selective GRU expert and the linear attention expert, and a gated MLP block for further learning and filtering complex user behaviors.
In the expert mixing block, the dense selective GRU module is employed to capture the long-term dependencies and local connections among the items, and selectively learn the sequential information. In addition, the linear attention expert is responsible for modeling item interactions from the user. By combining these two powerful expert modules, our GLINT-RU is capable of adaptively learning dependencies and relevant items from the sequence, which gains deeper insights into the complex user behaviors.
@@ 3.2. Item Embedding Layer
For sequential recommendation tasks, information on items interacted by users should be encoded to tensors through the embedding layer (Zhao et al., 2023). We denote the length of input user-item interactions as , and embedding size as . For a interaction sequence , the -th item can be projected into the representation by the following formulation:
@@
In traditional transformer-based models, positional embeddings are typically necessary because the attention mechanism lacks the inherent capability to encode temporal information (Kang and McAuley, 2018). Uniquely, in this paper we employ the GRU module to capture sequential behaviors, which has already taken temporal information into consideration. Therefore, we decide not to add the positional embedding layer into the framework.
@@ 3.3. Dense Selective GRU
Existing GRU cell learns sequential data by conveying information from preceding cells. Although this mechanism has the superiority of capturing the long-term dependencies in the sequence, it predominantly focuses on information from previous items, while potentially overlooking the valuable context information provided by adjacent items, which are often closely related in real-world applications. To address these challenges and extract local temporal features, we introduce the dense selective GRU as the core component of the GLINT-RU framework. This innovation aims to enhance the model’s flexibility and ability to capture complex long-/short-term patterns in user behavior. By implementing dense selective GRU, the computational complexity can be further reduced, and the recommendation accuracy of GRU-based framework can be substantially improved.
@@ 3.3.1. Dense GRU module
Therefore, to enable each GRU cell to be exposed to local temporal features of user behaviour, we introduce a temporal convolution layer, where adjacent item information is adaptively fused before being fed into GRU module:
where is the input tensor with feature dimensions, is the output of the linear layer, and is the output of the convolution operation with steps. and are weight matrix and bias, respectively. The size of the convolution kernel is set as . According to Eq.(1), the GRU cell generates the output based on the input and the hidden state at the previous time step:
where is the basic unit in Eq.(1) that processes a current hidden state. Then we generate new features by linearly combining the outputs of the GRU shown in Figure 1, capturing complex patterns of user preferences. These novel feature representations can be more effectively utilized for predicting behaviors and preferences of users. The feature-crossing process can be written as:
@@ 3.3.2. Selective Gate
To filter the hidden information of the GRU module, we design a selective gate where outputs of the feature crossing layer are selected based on the input state of the GRU. The selective gate weights are generated by a two tiny linear layers with SiLU activation function (Elfwing et al., 2018), and we use them to select useful hidden states and filter information:
@@ 3.4. Expert Mixing Block
Although a single GRU module can capture long-term dependencies for recommendation tasks, it lacks the ability to model global connections within the item sequence. By introducing the linear attention mechanism and mixing these two powerful experts, the computational complexity will be substantially reduced compared with transformer-based models, and more information about item connections and dependencies will be considered. Moreover, the two employed experts are parallel in our framework, which further improves the model efficiency.
In real applications, the conditions of the data might vary a lot. The GRU is naturally suited for sequential data, demonstrating effective performance in sequential recommendation tasks that exhibit strong temporal dependencies, while attention focuses on relevant items in the sequence dynamically. To adapt to complex data conditions, we attribute appropriate weights to the two experts by a mixing gate:
@@ 3.5. Gated MLP Block
Most existing frameworks, especially transformer-based models, utilize a two-layer feed forward network to capture the nonlinear relationships among features before giving predictions. To further enhance the performance of the model and augment useful features from the expert mixing block, we introduce the gated MLP block, which employs a gate mechanism again to deeply filter the information and generate item representations for predictions.
where is the output of expert mixing block, denotes the output of gated linear layer, represents the item representation, and are weight matrices, are bias. The recommendation scores are generated by item representations and embeddings, followed by final prediction probability of the -th item:
where is the representation of -th item. Finally, our GLINT-RU give top- items with the highest scores for recommendation tasks.

### v2

@@ Abstract.
Transformer-based models have gained significant traction in sequential recommender systems (SRSs) for their ability to capture user-item interactions effectively. However, these models often suffer from high computational costs and slow inference. Meanwhile, existing efficient SRS approaches struggle to embed high-quality semantic and positional information into latent representations. To tackle these challenges, this paper introduces GLINT-RU, a lightweight and efficient SRS leveraging a single-layer dense selective Gated Recurrent Units (GRU) module to accelerate inference. By incorporating a dense selective gate, GLINT-RU adaptively captures temporal dependencies and fine-grained positional information, generating high-quality latent representations. Additionally, a parallel mixing block infuses fine-grained positional features into user-item interactions, enhancing both recommendation quality and efficiency. Extensive experiments on three datasets demonstrate that GLINT-RU achieves superior prediction accuracy and inference speed, outperforming baselines based on RNNs, Transformers, MLPs, and SSMs. These results establish GLINT-RU as a powerful and efficient solution for SRSs. The implementation code is publicly available for reproducibility. 111https://github.com/szhang-cityu/GLINT-RU.
@@ 1. Introduction
In this era of data exploding, sequential recommender systems (SRSs) (Zhang et al., 2019; Yue et al., 2024; Liu et al., 2023a; Hidasi et al., 2015; Kang and McAuley, 2018; Liu et al., 2024a; Li et al., 2022; Liu et al., 2024b; Zhao et al., 2023a) have gained much attention in capturing users’ preferences within a large amount of sequential interaction data. GRU4Rec (Hidasi et al., 2015) as one of the earliest session-based recommendation models, employs stacked Gated Recurrent Units (GRU) for item-to-item recommendations. However, the RNN-based methods (Hidasi et al., 2015; Li et al., 2017) are fading from the recommendation realm due to their relatively lower accuracy. In recent years, transformer-based SRSs (Kang and McAuley, 2018; Sun et al., 2019) have become increasingly popular for the powerful multi-head attention mechanism (Zhao et al., 2023c; Vaswani et al., 2017). They exhibit remarkable ability in capturing sequential interactions and delivering accurate predictions (Li et al., 2023). However, despite their effectiveness, current transformer-based models suffer from substantial computational demands and extended inference time, which is caused by the dot product operation in the attention mechanism (Wang et al., 2020; Liu et al., 2023a; Zhou et al., 2023).
@@
To tackle the issue of the high computational cost of transformer-based SRSs, linear attention mechanisms are applied to reduce the computational complexity. For example, LinRec (Liu et al., 2023a) changes the order of dot product between query and key matrices by designing a special mapping, dramatically reducing the inference time. LightSAN (Fan et al., 2021) projects historical interactions into interest representations with shorter lengths, thereby reducing the computational complexity of transformers to a linear scale. MLP-based frameworks (Li et al., 2022; Zhou et al., 2022; Long et al., 2024) can also achieve fast inference speed and high performance by reshaping input sequence tensors (Gao et al., 2024). SSM-based models (Liu et al., 2024a; Yang et al., 2024) outperforms the existing attention mechanism by utilizing the efficient selective SSM (Gu and Dao, 2023), within which a structured state tensor is used to address long-range dependencies.
However, to achieve high performance, the transformer-based SRSs, even with linear attention mechanism, require deeply stacked transformers, which decreases the model efficiency (Zhang et al., 2024). MLP-based SRSs with sequence mixing layers might suffer from extended inference time when dealing with long sequences. In addition, MLP-based models struggle to capture fine-grained positional dependencies. SSM-based models (Liu et al., 2024a) may struggle to model effective semantic features into latent representations in long/short-term recommendation scenarios, as they might have difficulty learning important interactions. In this paper, we aim to further improve the efficiency and the accuracy of efficient models in various scenarios.
@@
To further reduce resource consumption, accelerate inference speed, and enhance the model performance, we propose a novel efficient SRS framework named Gated Lightweight IntelligeNT Recurrent Units (GLINT-RU). Considering that stacking hybrid architectures may lead to a deeper model, which will cause a significant decrease in inference speed. and the traditional GRU module lacks the ability to adaptively adjust the output and filter out unimportant interactions. To tackle these challenges, we employ various gate mechanisms in appropriate positions to fully perceive the data environment and filter information automatically (Zhang et al., 2024; Chang et al., 2023; De et al., 2024). We introduce an expert mixing block that captures the item dependencies via GRU and utilizes linear attention to capture the global interaction information between users and items (Jacobs et al., 1991; Masoudnia and Ebrahimpour, 2014). This strategy not only improves the inference speed due to the linear computational complexity of paralleled GRU and linear attention mechanism but also enhances the context information. Additionally, we implement a dense selective GRU, which selects the output of the GRU adaptively and considers the connections among adjacent items. It leverages the gate mechanism to select crossed channels and extracts short-duration patterns to refine the model’s understanding of user behavior dynamics. Moreover, a gated MLP block is utilized to select the outputs of the expert mixing block, deeply filtering the information based on the data environment.
We summarize the major contributions of our work as follows:
@@
In this paper, we introduce GLINT-RU, a novel and lightweight SRS that achieves remarkable inference speed only requiring a single layer. It is an advanced model that captures complex semantic features and fine-grained positional representations using an expert mixing mechanism, which substantially improves the performance of GLINT-RU over existing efficient SRS models.
@@
We introduce a dense selective GRU module, which not only incorporates connections between adjacent items but also empowers the model with the capability to selectively learn long-term dependencies. The integration of this advanced GRU module into the model markedly elevates its performance, establishing a new standard for efficient recommender systems.
We conduct extensive experiments to verify the efficiency and effectiveness of GLINT-RU on various datasets and parameter settings. Our GLINT-RU improves the training and inference speed significantly and stably exhibits high performance.
@@ 2.1. Problem Statement
For a sequential recommendation task, we have a set of users who have historical interactions with a set of items . Among these users, the -th user has a preferred item sequence denoted as , , where is the length of the item list that the -th user interacts with. Our goal is to design an efficient framework and predict the rating score of the next item based on the historical interactions .
@@ 2.2. Gated Recurrent Units
As an essential part of GLINT-RU, the GRU (Chung et al., 2014) module contributes to the recommendation task by capturing the dependencies among the items and dynamically adjust its memory content (Dey and Salem, 2017; Shen et al., 2018). The update mechanism of a GRU cell is formulated as follows:
@@
The GRU (Gated Recurrent Unit) module, equipped with update and reset gates in sequential GRU Cells, is adept at capturing the relationships among the items throughout the sequence while maintaining a relatively low computational complexity. However, the sequential information in GRU cannot be interacted with, and each hidden state is primarily encoded from preceding elements, which restricts the representational capacity of GRU-based recommendation models, particularly in capturing complex item dependencies across the entire sequence.
@@ 2.3. Linear Attention Mechanism
The attention mechanism as a powerful core of the transformer structure, exhibits performance in learning sequence interactions in recommendation tasks. However, the high computational cost of the dot product between query matrix and key matrix substantially lower the inference speed of transformer-based SRSs especially when the sequence length is much larger than hidden size (Shen et al., 2021; Wang et al., 2020). To tackle this issue, the linear attention mechanism (Liu et al., 2023a) designs a special mapping function to change the order of the dot product and reduce the computational complexity to . The linear attention mechanism can be written as:
where and are row-wise and column-wise normalization mappings, respectively, are learnable query, key and value matrices and is the output attention score. This approach mitigates the issue that the softmax layer concentrates on the scores of merely a few positions, enlarging the information capacity of the attention mechanism (Liu et al., 2023a). By implementing linear attention, our GLINT-RU framework is capable of learning interactions between items in long sequences.
@@ 3. METHODOLOGY
In this section, we will introduce the overall framework and its components that effectively capture semantic features and positional information, followed by the complexity analysis of GLINT-RU.
@@ 3.1. Framework Overview
Many existing recommendation frameworks depend on transformer structure (Sun et al., 2019; Kang and McAuley, 2018; Liu et al., 2023a), which incurs substantial computational overhead and low inference speed. Restricted by the large computational complexity of stacked transformers, linear attention-based models have approached a plateau in terms of minimizing inference time and resource consumption. Uniquely, we propose an advanced recommendation framework that integrates the linear attention mechanism and efficient dense selective GRU module, which further reduces the computational cost compared with stacked linear transformers and SSM-based models. Additionally, this dense selective GRU module also enables our framework to understand both semantic features and dependencies from item sequences, and substantially reduce the computational cost and inference time.
Figure 1.(a) shows the structure of our GLINT-RU. As is shown in Figure 1.(b)-(d), GLINT-RU integrates an expert mixing block for mixing sequential information from the dense selective GRU expert and the linear attention expert, and a gated MLP block for further learning and filtering complex user behaviors.
In the expert mixing block, the dense selective GRU module is employed to capture the long/short-term item-wise dependencies, and selectively learn the sequential information. In addition, the linear attention expert is responsible for modeling item interactions from the user. By combining these two powerful expert modules, our GLINT-RU is capable of adaptively learning both temporal and semantic item features from the sequence, which performs fine-grained modeling of complex user behaviors.
@@ 3.2. Item Embedding Layer
For sequential recommendation tasks, information on items interacted by users should be encoded to tensors through the embedding layer (Zhao et al., 2023b). We denote the length of input user-item interactions as , and embedding size as . For a interaction sequence , the -th item can be projected into the representation by the following formulation:
@@
In traditional transformer-based models, positional embeddings are typically necessary because the attention mechanism lacks the inherent capability to encode temporal information (Kang and McAuley, 2018). Uniquely, in this paper we employ the GRU module to model temporal item dependencies, which generates fine-grained representations with positional information for the items. Therefore, we decide not to add the positional embedding layer into the framework.
@@ 3.3. Dense Selective GRU
Existing GRU cell learns sequential data by conveying information from preceding cells. Although this mechanism has the superiority of capturing the long-term dependencies in the sequence, it predominantly focuses on information from previous items, while potentially overlooking the valuable context information provided by adjacent items, which are often closely related in real-world applications. To address these challenges, we introduce dense selective GRU shown in Figure 1.(b) as the core component of GLINT-TU. This innovation extracts local temporal features and generates fine-grained positional information using the update mechanism in GRU module. By implementing dense selective GRU, the computational complexity can be further reduced, and the recommendation accuracy of the GRU-based framework can be substantially improved.
@@ 3.3.1. Dense GRU module
Therefore, to enable each GRU cell to aggregate local temporal features of user behavior, we introduce a temporal convolution layer , where adjacent item information is adaptively fused before being fed into the GRU module:
where is the input tensor with feature dimensions, and is the output of the convolution operation with steps. and are weight matrix and bias, respectively. The size of the convolution kernel is set as . According to Eq.(1), the output of the can be divided into a latent item representation and a fine-grained positional representation learned by historical hidden states:
where is the reset gate at the -th time step. It is noteworthy that each positional representation is generated by aggregated historical hidden states with varied update intensities. Then we project the hidden states into a latent space using the channel crossing layer, which can be written as:
@@ 3.3.2. Selective Gate
To filter the hidden information of the GRU module, we design a selective gate where outputs of the feature crossing layer are selected based on the input state of the GRU. The selective gate weights are generated by two tiny linear layers with SiLU activation function (Elfwing et al., 2018), and we use them to select useful hidden states and filter information:
@@ 3.4. Expert Mixing Block
Existing efficient SRSs lack the ability to exploit semantic features, fine-grained positional information, and dependencies simultaneously. As is shown in Figure 1.(c), by introducing the linear attention mechanism and mixing these two powerful experts, the computational complexity will be substantially reduced compared with transformer-based models, and more high-quality item representations can be generated. Moreover, the two employed experts are parallel in our framework, which further improves the model’s efficiency. The expert mixing mechanism is shown in Figure 2.
In real applications, the conditions of the data vary a lot. The GRU is naturally suited for sequential data, demonstrating effectiveness in sequential recommendation tasks that exhibit strong temporal dependencies, while attention focuses on relevant items in the sequence dynamically. To adapt to complex data conditions, we attribute appropriate weights to the two experts by a mixing gate:
@@ 3.5. Gated MLP Block
Most existing efficient SRSs, utilize two-layer feed forward networks to capture the nonlinear relationships among features before giving predictions. To further enhance the performance of the model and augment useful features from the expert mixing block, we introduce the gated MLP block shown in Figure 1.(d), which employs a gate mechanism again to deeply filter the information and generate item representations for predictions.
where is the output of expert mixing block, denotes the output of gated linear layer, represents the item representation, and are weight matrices, are bias. The recommendation scores are generated by item representations and embeddings, followed by the prediction score of the -th item:
where is the representation of -th item. Loss function, model training methods and the algorithm are displayed in Appendix A.